Use the following information from questions 6-12. In both 2018 and 2019 you expect to receive a W2 for $200,000; $65,000 will be already withheld on your W2 for federal income taxes and your income tax bracket is 32%. The short term capital gains on investments (<1 year) are taxed like income, long term capital gains on investments (>=1 year) are taxed at 15%, and dividends are taxed at 22%. Besides your W2 income, the only income you have to report is from one stock on the 1099-B issued by your broker: On January 1t, 2018, you purchased 100 shares of stock at $50 per share; you based your investment decision on publically available sources which speculate the shares price will increase dramatically at some point over the next two years July 1s, 2018, is the ex-dividend date for the stock - that morning you received a special onetime cash dividend of $2 per share and the stock closed at $48 per share On November 29h, 2018, the company announced great news about a clinical drug trial and the share price begins to rise On November 30th, 2018, you purchased 20 shares of stock at $80 per share On December 1st, 2018, there was a 2 for 1 stock split On December 2nd, 2018, you sold 50 shares of stock at $100 per share On January 2st, 2019, you sold 70 shares at $90 per share You do not receive any dividends in 2019 On December 31s, 2019, you sell any remaining shares at $30 per share All stock prices quoted above are at market close on the day referenced Assume shares that were purchased first were sold first (i.e. FIFO) Assume there are no transaction costs

After the sale of the stock on December 2nd, 2018, your cost basis per share is closest to:

a. $27.50 b. $28.16 c. $29.29 d. $50.83 e. $52.50
Answer #1

Answer: Option A: $27.50

Total share value on December 1st =100*50+20*80=$7600

and total number of shares=100+20=120

after 2 for 1 stock split new number shares= 120*2=240

So Cost basis per share = total share value /New number of shares=7600/240=$27.5